# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ 3.3 Pruning-Based Mitigation Conceals but Does Not Erase Memorization
Table 1: Pruning-based mitigation of memorization is vulnerable to adversarial embeddings. Without any mitigation technique applied (1st row), the generated images clearly indicate data replication. Searching for adversarial embeddings on non-memorized prompts (2nd row) does not lead to clear replication. After localizing and pruning weights with NeMo (3rd row) or Wanda (4th row), data replication appears effectively prevented. However, identifying adversarial embeddings with Dori—indicated by [Uncaptioned image]—reveals that embeddings capable of triggering data replication may persist, even after pruning. Finally, our adversarial fine-tuning (5th row) successfully removes memorized content and prevents data replication. N/A denotes not applicable.
Setting Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ $\boldsymbol{\downarrow}$FID
No Mitigation Verbatim $\bm{0.90}\pm 0.04$ N/A $1.00\pm 0.00$ $0.33\pm 0.01$ $14.44$
Non-Memorized Prompts None $0.17\pm 0.05$ N/A $0.35\pm 0.06$ $0.35\pm 0.02$ $14.44$
Non-Memorized Prompts + [Uncaptioned image] None $\bm{0.48}\pm 0.06$ N/A $0.67\pm 0.07$ $0.32\pm 0.02$
NeMo [14] Verbatim $0.33\pm 0.18$ $0.40\pm 0.21$ $0.46\pm 0.13$ $0.34\pm 0.02$ $15.16$
NeMo + [Uncaptioned image] Verbatim $\bm{0.91}\pm 0.03$ $0.97\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$
Wanda [3] Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $0.34\pm 0.02$ $16.86$
Wanda + [Uncaptioned image] Verbatim $\bm{0.76}\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$
Adv. Fine-Tuning (Ours) Verbatim $0.15\pm 0.07$ $0.15\pm 0.07$ $0.35\pm 0.08$ $0.33\pm 0.01$ $13.61$
Adv. Fine-Tuning (Ours) + [Uncaptioned image] Verbatim $\bm{0.36}\pm 0.14$ $0.54\pm 0.10$ $0.54\pm 0.10$ $0.30\pm 0.02$
@@
We begin by showing that NeMo and Wanda prevent data replication only in the text space—where users input prompts—but do not fully remove the memorized content from the DM. As shown in the first row of Tab. 1, verbatim memorized prompts trigger replication in the unmodified DM. In contrast, image generations for non-memorized training prompts (2nd row) show no signs of memorization under the SSCD score. Even when applying our adversarial embedding optimization, indicated by [Uncaptioned image] in the table, the resulting metrics suggest no close data replication. This finding is particularly important for the validity of our investigations, as it confirms that our adversarial embedding optimization method specifically targets memorized content and does not falsely indicate memorization for non-memorized data in the current setting. We explore adversarial embeddings in the context of non-memorized content more closely in Sec. D.1.
Applying NeMo (third row) or Wanda (fourth row), which identify and prune weights associated with memorization in the cross-attention value and output fully-connected layers, respectively, substantially reduces data replication as reflected by low SSCD scores. At first glance, both methods appear effective at mitigating data replication, as also visualized in Fig. 1 ( 2). However, blocking replication from the original prompts does not imply that the memorized content has been removed from the model. To probe this, we employ adversarial embedding optimization (Sec. 3.1) to explore the neighborhood around memorized prompt embeddings.
@@
Despite pruning, we find that adversarial embeddings can still trigger replication, as shown in rows marked with [Uncaptioned image] and Fig. 1 ( 3), particularly for VM prompts. These results suggest that pruning-based methods like NeMo and Wanda primarily conceal memorization rather than eliminate it, preventing replication via the text space but leaving the memorized content internally intact. We also conduct a sensitivity analysis (Sec. E.2) on the steps required to yield adversarial embeddings, finding that in most cases, significantly fewer than 50 steps are already sufficient to identify embeddings that circumvent the mitigation methods. For TM results, also reported in Sec. E.2, we observe increased replication, yet SSCD-based metrics fail to correctly quantify this type of replication due to their non-semantic variations in generated images.
Based on the presented results, one might argue that the number of pruned weights simply needs to be increased to fully mitigate memorization. However, we show that this is not the case. For Wanda, we find that increasing the strength of pruning, i.e., removing more weights, successfully defends against adversarial embeddings, but at a cost of significant damage to generation quality, as shown in Sec. E.7. Specifically, to completely remove a single image, one needs to prune about 10% of the weights in the output fully-connected layers with Wanda. At this scale, the DM loses its capability to reliably generate concepts related to the memorized image, even from non-memorized prompts.
Since NeMo lacks a parameter to directly control the pruning budget, we instead adopt an iterative approach. In Sec. E.6, we combine NeMo’s pruning method with adversarial embedding search, yielding a more robust and lasting defense. Specifically, we first identify memorization weights with NeMo for original memorized text embeddings, and prune them from the model. We then perform our adversarial embedding optimization to obtain a new embedding that elicits data replication, identify a new set of weights for this adversarial embedding, and prune them, while keeping the previously pruned weights inactive. This process is repeated, expanding the pruned set with each iteration. Despite this effort, we find that data replication persists, and the model’s utility degrades substantially as more weights are removed.
Given the fragility of pruning-based approaches, we modify our adversarial search technique (Sec. 3.1) by initializing the optimized embeddings by sampling from $\mathcal{N}(\bm{0},\bm{I})$, instead of using the memorized embeddings as initialization, and repeating the experiments from this section. The results, presented in Sec. E.3, closely match the results in Tab. 1, indicating that memorization triggers are not local. We explore this phenomenon in greater depth in the next section.
@@ 4 The Illusion of Memorization Locality
Our finding that pruning only conceals memorization (Sec. 3.3), requiring severe degradation for full removal, contradicts the locality assumption, i.e., that only a small set of weights is responsible for memorizing a specific training sample. This motivates us to investigate its validity more closely. We examine the text embedding space and internal DM activations, revealing that evenly distributed adversarial triggers elicit distinct internal DM activation patterns, resulting in low overlap among weights identified by pruning. This evidence suggests the locality assumption to be incorrect.
@@ 4.1 Data Replication Triggers are Not Localized in Text Embedding Space
Existing memorization mitigation techniques operate solely on the text embedding of the memorized prompt, assuming that disentangling the generation process from the prompt’s embedding is sufficient to prevent replication. In the previous section, we focused our adversarial embedding search on local subspaces around the embedding of the memorized training prompt. Assuming that all adversarial embeddings for a specific memorized image indeed share the same local subspace, a natural mitigation would be to construct a region around each memorized embedding, e.g., an $\epsilon$-ball, and map all points within this region back to its center.
However, our findings contradict this assumption: adversarial embeddings for a single memorized image can be found evenly distributed in the text embedding space. To demonstrate this fact, we craft a set of 100 adversarial embeddings, denoted by $\bm{Y}_{adv}$, for the same memorized image. Instead of initializing the embeddings with the memorized prompt, initialization is performed at random positions, i.e., for each ${\bm{y}}_{adv}\in\bm{Y}_{adv}$ we have ${\bm{y}}^{\smash{(0)}}_{\text{adv}}\sim\mathcal{N}(\bm{0},\bm{I})$.
@@
After optimizing each embedding for 50 steps using the procedure described in Sec. 3.1, we generate the corresponding images and assess their similarity to the memorized training sample. Across all runs, the generated images consistently yield $\bm{\text{SSCD}_{\text{Orig}}}$ scores above $0.7$, strongly indicating successful replication. Yet, the optimized embeddings do not collapse to a single point but are scattered in the embedding space, as visualized by the t-SNE [42] plot in Fig. 2. We repeat the experiment by initializing ${\bm{y}}^{\smash{(0)}}_{\text{adv}}$ with embeddings of 100 randomly selected, non-memorized prompts. Results from this experiment, presented in Fig. 6 (Sec. G.1), draw a similar picture of evenly distributed replication triggers. Both results clearly demonstrate that data replication can be triggered virtually from all over the embedding space, taking away the illusion of local memorization triggers.
Figure 2: Data replication triggers are scattered in the embedding space. We show that the adversarial embeddings are distributed similarly to the randomly initialized embeddings, thus refuting the input space locality of data replication triggers.
We continue by providing more quantitative support for our findings by computing the pairwise distances in the set of randomly initialized embeddings and repeating the same for the set of optimized adversarial embeddings. Our findings, also visualized in Fig. 7 (Sec. G.1), show that the optimized embeddings of $\bm{Y}_{adv}$ are even more spread out than their random initializations. This result highlights that the adversarial embeddings are widely scattered and do not converge to a single region.
In the embedding space, we find no traces of memorization locality in the sense that only local regions can trigger replication of a specific memorized image—an assumption implicitly made by existing pruning-based mitigation methods.
Successful mitigation of memorization should, therefore, drop this locality assumption and account for a significantly broader range of adversarial embeddings beyond the space around the memorized training prompt embedding.
@@ 4.2 Images are Not Memorized in a Subset of Weights
Building on our observation of high diversity among adversarial embeddings, we now turn our attention to internal model activations. For fixed input noise, we anticipate these activations will vary depending on the text embedding provided to the DM. This analysis is crucial, because pruning methods like Wanda and NeMo rely heavily on internal activations—treated as per-weight metrics—to identify weights contributing to data replication as candidates for pruning. Consequently, if different adversarial embeddings leading to the same output yield distinct activation patterns, we would expect Wanda and NeMo to prune different sets of weights. Inconsistency in the resulting pruned sets would suggest a lack of empirical support for the locality assumption regarding memorization.

### v2

@@ 3.3 Pruning-Based Mitigation Conceals but Does Not Erase Memorization
Table 1: Pruning-based mitigation of memorization is vulnerable to adversarial embeddings. Without any mitigation technique applied (1st row), the generated images clearly indicate data replication. Searching for adversarial embeddings on non-memorized prompts (2nd row) does not lead to clear replication. After localizing and pruning weights with NeMo (3rd row) or Wanda (4th row), data replication appears effectively prevented. However, identifying adversarial embeddings with Dori (indicated by [Uncaptioned image]) reveals that embeddings capable of triggering data replication may persist, even after pruning.
[…]
No Mitigation Verbatim $\bm{0.90}\pm 0.04$ N/A $0.33\pm 0.01$ 0.98 $14.44$
Non-Memorized Prompts None $0.17\pm 0.05$ N/A $35\pm 0.02 0.0$0 $14.44$
[…]
NeMo (Hintersdorf et al., 2024) Verbatim $0.33\pm 0.18$ $0.40\pm 0.21$ $0.34\pm 0.02$ 0.20 $15.16$
NeMo + [Uncaptioned image] Verbatim $\bm{0.91}\pm 0.03$ $0.97\pm 0.02$ $0.33\pm 0.02$ $\bm{0.99}$
Wanda (Chavhan et al., 2024) Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $34\pm 0.02 0.0$0 $16.86$
Wanda + [Uncaptioned image] Verbatim $\bm{0.76}\pm 0.05$ $0.82\pm 0.05$ $0.32\pm 0.01$ $\bm{0.72}$
@@
Our analysis in Table˜1 highlights that NeMo and Wanda prevent training data replication only in the text space, i.e., breaking the mapping from the prompt to the corresponding memorized image, but do not remove the memorized images from the DM. As shown in the first row of Table˜1, verbatim memorized prompts trigger the replication of the memorized training images in the original DM (No Mitigation). In contrast, image generations for non-memorized training prompts (2nd row) show no signs of memorization under the SSCD score. Even when applying Dori, our adversarial embedding optimization, indicated by [Uncaptioned image] in the table, the resulting metrics suggest no close data replication for non-memorized prompts. This finding is particularly important for the validity of our investigations, as it confirms that our adversarial embedding optimization method specifically targets memorized content and does not falsely report memorization for non-memorized data. We explore adversarial embeddings in the context of non-memorized data points more closely in Section˜F.1.
@@
Applying NeMo (third row) or Wanda (fourth row), respectively, substantially reduces training data replication as reflected by lowered SSCD scores in contrast to the original DM. At first glance, both methods appear effective at mitigating the replication of memorized data points, as also visualized in Fig.˜1 ( 2 ). However, blocking replication from the original prompts does not imply that the memorized individual images have been removed from the model, as shown in rows marked with [Uncaptioned image] and Fig.˜1 ( 3 ). These results suggest that pruning-based methods like NeMo and Wanda primarily conceal memorization rather than eliminate it. Also for TM results, reported in Section˜H.2, we observe increased replication of memorized training data, yet SSCD-based metrics fail to correctly quantify this type of replication due to their non-semantic variations in generated images. Overall, these results suggest that pruning-based memorization mitigations prevent the replication of memorized images via the text space but leave the memorized images intact internally in the DM.
Ablations. We also conduct a sensitivity analysis (Section˜H.2) on the number of steps required to yield adversarial embeddings, finding that in most cases, significantly fewer than 50 steps are already sufficient to identify embeddings that circumvent the mitigation methods. We also experiment with increasing the strength of NeMo and Wanda to test if they become resilient to Dori. For the former, we iteratively increase the set of pruned weights based on new adversarial embeddings (see Section˜H.6), and for the latter, we simply prune 10% of weights, instead of 1% (see Section˜H.5). In these settings, Wanda successfully removes memorization but at substantial damage to the DM’s generative capabilities (see Section˜H.7), while NeMo remains non-robust to Dori.
@@ 4 The Illusion of Memorization Locality
Our analysis in the previous section revealed that pruning-based memorization mitigation methods fail to remove memorized images from DMs, at least without compromising overall generation quality. This suggests that their underlying assumption of localized memorization is flawed. In this section, we provide multiple lines of evidence that memorization in DMs is indeed not inherently local. Specifically, in Section˜4.1, we show that replication triggers for memorized images are widely distributed in the text embedding space, in Section˜4.2, we demonstrate that distinct embeddings producing the same memorized training image can induce divergent model activations, and in Section˜4.3, we find that different pruning methods identify inconsistent sets of memorization-related weights for the same image. Finally, in Section˜4.4 we demonstrate that abandoning the locality assumption enables more robust mitigation through adversarial fine-tuning, paving the way towards novel methods that truly mitigate memorization of individual data points in DMs.
@@ 4.1 Data Replication Triggers are Not Localized in Text Embedding Space
Figure 2: Data replication triggers are widely and uniformly scattered in the text embedding space.
In the previous section, we demonstrated that pruning-based memorization mitigations (NeMo and Wanda) can be broken by optimizing adversarial embeddings initialized near the original training prompt. Here, we show that triggers in the embedding space do not need to lie close to a memorized image’s prompt: even adversarial embeddings initialized at random, distant positions in the text embedding space can still reliably trigger replication of memorized images.
To illustrate this, we construct a set of 100 adversarial embeddings, $\bm{Y}_{adv}$, for a single memorized image, where each embedding is randomly initialized: ${\bm{y}}^{(0)}_{\text{adv}}\sim\mathcal{N}(\bm{0},\bm{I})$. After optimizing each embedding for 50 steps in accordance with the procedure described in Section˜3.1, we show that every run produces a generated image highly similar to the memorized sample (all with $\bm{\text{SSCD}_{\text{Orig}}}$ scores clearly exceeding $0.7$). Despite their consistent success at replication, the optimized adversarial embeddings remain widely dispersed in the embedding space and retain a distribution similar to their random initializations, as visualized by the t-SNE (van der Maaten & Hinton, 2008) plot in Fig.˜2. This result further refutes the assumption of input space locality.
@@
We repeat the experiment by initializing ${\bm{y}}^{\smash{(0)}}_{\text{adv}}$ with embeddings of 100 randomly selected, non-memorized prompts. Results from this experiment, presented in Fig.˜8 (see Section˜J.1), draw a similar picture of evenly distributed replication triggers. Both results clearly demonstrate that replication of memorized images can be triggered virtually from all over the embedding space, taking away the illusion of local memorization triggers.
To quantify our observations further, we compute the pairwise distances both among all the initial random embeddings and among all the optimized adversarial embeddings. As shown in Fig.˜9 (Section˜J.1), the optimized embeddings are in fact even more dispersed than their initializations, confirming that successful replication triggers are widely scattered rather than localized. These results demonstrate that the assumption of input locality, implicit in the pruning-based mitigation methods, is unfounded. Effective memorization mitigation must therefore address potential triggers distributed throughout the embedding space, not just those near the training prompt.
4.2 Different Triggers, Different Activations, the same Image
Next, we examine internal model activations to assess whether replication triggers for memorized images are also dispersed at the level of network activity. For fixed input noise, we expect that activations should vary depending on the guiding text embedding. This analysis is essential, as pruning methods like Wanda and NeMo select candidate weights for removal based on activation patterns, treating these as per-weight importance metrics. If different adversarial embeddings that trigger the same image yield distinct activations, these methods may prune inconsistent sets of weights, undermining the assumption of locality in memorization further.
To quantify activation variability, we introduce a discrepancy metric, defined as the mean pairwise $\ell_{2}$-distance between activations in a given layer across different input embeddings during the initial denoising step. To ensure a fair comparison, the input noise is fixed while only the guiding adversarial embedding is varied. The exact formulation is provided in Eq.˜6 (see Section˜J.2). We compute discrepancies for two embedding sets: $\bm{Y}_{adv}$, comprising 100 adversarial embeddings for a single memorized image (from Section˜4.1), and $\bm{Y}_{\mathit{mem}}$, containing 100 text embeddings of randomly selected prompts associated with different memorized images. Since replicating different images should produce more varied activations, we expect higher discrepancies for $\bm{Y}_{\mathit{mem}}$ and lower, more consistent values for $\bm{Y}_{adv}$.
Figure 3: Diverse activations refute locality. Although adversarial embeddings trigger the same image, their activations exhibit high discrepancy.
We analyze activations from the layers where NeMo and Wanda operate, specifically, the value layers of cross-attention modules for NeMo and the second linear layers of the transformer blocks’ two-layer feed-forward networks for Wanda. In total, we compute activations for seven cross-attention modules, indexed from 1 to 7, spanning the three Down blocks (indices 1 to 6, each block has two modules) and the Mid block of the DM’s U-Net (Ronneberger et al., 2015), following the setup of NeMo.
Surprisingly, as shown in Fig.˜3, the discrepancy among activations for adversarial embeddings in $\bm{Y}_{adv}$ is comparable to that for randomly selected memorized prompts in $\bm{Y}_{\mathit{mem}}$. This finding indicates that different adversarial embeddings, even when generating the same image, cause distinct activation pattern, contradicting the expectation that a common output implies similar activations. While Wanda shows slightly reduced discrepancy for $\bm{Y}_{adv}$, the variability remains substantial, suggesting each adversarial embedding invokes a unique activation pattern. Extending this analysis to all U-Net layers (see Fig.˜10 in Section˜J.3) yields similar results, providing further evidence against the locality assumption in memorization.
@@ 4.3 Images are Not Memorized in a Subset of Weights
While high discrepancy scores suggest that different subsets of weights contribute to data replication, we further assess the consistency of pruning-based mitigation methods across adversarial embeddings for the same image. Since NeMo and Wanda select weights for pruning based on activation patterns, we expect the identified sets of weights to vary with different adversarial triggers. To quantify this, we define a weight agreement metric as the intersection over union of weights identified for pruning by NeMo or Wanda between two adversarial embeddings, averaged across all pairs. Higher values indicate greater overlap of the identified weights. The metric’s formal definition is given in Eq.˜7 (see Section˜J.2). Weight agreement is set to 1, representing a perfect overlap, if no weights are selected for either embedding in a given layer.
Figure 4: Locality fails in the model’s weights. Large activation discrepancy (Fig.˜3) results in low weight agreement, further undermining the idea that weights responsible for replicating a memorized image can be pinpointed and pruned.
As shown in Fig.˜4, the overlap in identified weights for adversarial prompts in $\bm{Y}_{adv}$ is limited: Wanda ’s agreement remains below 0.6 for most layers, while NeMo exceeds 0.8 except in the first layer, where the overlap drops to about 0.6, similar to results for the set of diverse memorized prompts $\bm{Y}_{\mathit{mem}}$. Notably, this high agreement in deeper layers is mainly due to NeMo attributing most memorization-related weights to the first layer, as further visualized in Fig.˜11 (left) in Section˜J.2. Despite appearing more stable, iterative pruning experiments (see Section˜H.6) reveal that NeMo still identifies different weights for different adversarial embeddings after previous weights are pruned.
Crucially, the weight agreement among adversarial embeddings in $\bm{Y}_{adv}$ is comparable to that among distinct memorized prompts in $\bm{Y}_{\mathit{mem}}$, reinforcing the finding that both pruning-based approaches struggle to consistently localize the weights responsible for memorization. This further challenges the locality assumption and underscores the limitations of such memorization mitigation strategies.
4.4 Abandoning the Locality Assumption Yields Robust Mitigation Method
The consistent evidence against memorization locality from the previous sections suggests that memorization removal in DMs cannot be achieved by targeting just a subset of weights. Instead, our findings point to the need for mitigation strategies that operate at the level of the entire model. To verify this, we evaluate a simple but powerful method: adversarial fine-tuning, where all model parameters are adjusted to remove memorized content.
Approach. We employ Dori (Section˜3.1) to generate multiple adversarial embeddings for each memorized image and fine-tune the DM in an adversarial manner, inspired by adversarial training (Szegedy et al., 2014; Goodfellow et al., 2014; Madry et al., 2018). During training, adversarial embeddings should produce images that are semantically close to the memorized samples, but not exact replicas. To facilitate this, we first craft a set of surrogate images $\widetilde{{\bm{x}}}$, obtained by prompting the DM with memorized prompts under a pruning-based mitigation. These surrogate images resemble the memorized samples in content and style, yet differ in their details, ensuring they are no replications (see Section˜I.2 for visual examples). In addition, at each fine-tuning step, we iteratively collect a set of adversarial embeddings ${\bm{y}}_{adv}$ that trigger replication of the memorized images. We then fine-tune the DM on a mix of these surrogate images and adversarial embeddings with a loss defined as:
$\mathcal{L}_{Adv}(\widetilde{{\bm{x}}}_{0},\bm{\epsilon},{\bm{y}}_{adv},t,\bm{\theta})=\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}\left(\widetilde{{\bm{x}}}_{t},t,{\bm{y}}_{adv}\right)\|_{2}^{2}\text{.}$ (2)
The loss encourages the model to avoid replicating the memorized training image triggered by the adversarial embedding. Instead, it guides generation toward the surrogate images, which preserve the semantic content without being exact copies. Using a diverse set of surrogate images ensures that no new memorized image is inadvertently introduced into the DM. In addition, we use the standard diffusion loss $\mathcal{L}_{\text{DM}}$ defined in Eq.˜3, to train on non-memorized image-captions pairs from the LAION (Schuhmann et al., 2022) dataset, to preserve the model’s general utility. The final optimization loss is $\mathcal{L}=\mathcal{L}_{\text{DM}}+\mathcal{L}_{Adv}$. An algorithmic description of our adversarial fine-tuning method, as well as the full setup, is provided in Appx.˜I.
Results. We find that our adversarial fine-tuning procedure quickly removes memorized content. Table˜2 (bottom row) presents the evaluation results after fine-tuning for five epochs. The results show that adversarial embeddings can no longer trigger data replication (except for a single, highly duplicated training sample), demonstrating a permanent mitigation relative to pruning-based methods. At the same time, the model’s utility is preserved: the FID score improves from 14.44 to 13.61 after fine-tuning, suggesting no harm to the image quality, as extended results in Tab.˜3 (see Appx.˜G) show. We provide comprehensive analyses of the parameters and performance of the method in Appx.˜I. These results indicate that even a single fine-tuning epoch substantially reduces memorization and can prevent data replication in most cases. We also experimented with fine-tuning the DM with LoRA adapters (Hu et al., 2021), but found it unsuccessful in mitigating memorization, further underscoring that effective memorization mitigation requires global model adjustments.
Table 2: Dori [Uncaptioned image] against fine-tuning removal.